Factor (x-2)^5 -9(x-2)^3 =0
Answer by DrBeeee(684)


Given: (x-2)^5 - 9(x-2)^3 = 0
Factor out the GCF yields,
[(x-2)^3][(x-2)^2 - 9] = 0
Simplifying the second factor gives
[(x-2)*(x-2) - 9]
[x^2 - 2x - 2x + 4 - 9]
[x^2 -4x -5] which factors into
(x + 1)(x - 5)
The final factorization is
(x+1)(x-5)(x-2)^3 = 0